To be perfectly honest to you...I suck at word problems.
Here is the problem; Find two numbers such that their sum 10 and their product is 22.

Answer by checkley71(8403)

IF YOU EXPAND 5+/-SQRT3 YOU GET 5+1.73=6.73 & 5-SQRT3=3.27. WE JUST USED TWO DIFFERENT WAYS OF EXPRESSING THE ANSWER.

X+Y=10 THE X=10-Y & X*Y=22 THEN SUBSTITUTING 10-Y FOR X WE GET (10-Y)Y=22 OR
10Y-Y^2=22 OR Y^2-10Y+22=0 OR (Y-6.73)(Y-3.27)=0 OR Y=6.73 & Y=3.27
PROOF 6.73+3.27=10 OR 10=10 & 6.73*3.27=22 OR 22=22
I BELIEVE YOU MIGHT HAVE COPIED DOWN THE WRONG FIGURES. AS THESE NUMBERS DON'T LOOK LIKE A VALID SOLUTION.
PRODUCT=21 MAKES MORE SENSE THEN THE NUMBERS WOULD BE 3 & 7.